Calculate the ratio of pH of a solution containing 1 mole of CH3COONa + 1 mole of HCI per litre and of other solution containing 1 mole CH3COONa + 1 mole of acetic acid per litre. detailed solution plz

Grade:12th pass

2 Answers

Mr David
6 years ago
H+^ + CH3COO-^=CH3COOH1. 1. 0 1pH=pka+log(1÷1) solve to getpH (1)=pkaSimilarly pH2=pka+ log (conj base÷acid) which is also 1;pH2=pka Ratio =1:1 . I think the data of HCl is wrong .It is 0.5moles👍👍👍
Mudit Dharma
5 years ago
First 1 there is no buffer since only HCl gives 1 mole of H+ which gives pH -1.      2nd One there is buffer where pH=pKa as log 1/1 =0.  pKa =4.75 as Ka=1.75*10^-5 you get the ration as -2.01